1/a-1 + 3/3a-1 = 0
Answer by unreal4evr(11)

3/3a-1 = -(1/a-1)
3(a-1)/3a-1 = -1
3a-3/3a-1 = -1
3a/3a-1 - 3/3a-1 = -1
(3a-1)*(3a/3a-1 - 3/3a-1) = (3a-1) * -1
3a - 3 = -3a + 1
6a - 3 = 1
6a = 4
a = 4/6


//Check
1/(4/6)-1 + 3/3(4/6)-1 = 0
-3 + 3 = 0
0 = 0